A wave is modeled by the wave function y(x,t) = (0.66m) sin[(2π/7.58m)(x-(12.01m/s)t)]. What is the period of the wave?

Round your answer to 2 decimal places


1
Expert's answer
2021-05-03T10:33:28-0400
"T=\frac{2\pi}{\omega}\\T=\frac{2\pi(7.58)}{2\pi (12.01)}=0.63\ s"